# VideoHandles: Editing 3D Object Compositions in Videos Using Video Generative Priors

### v1

@@ Abstract
Generative methods for image and video editing use generative models as priors to perform edits despite incomplete information, such as changing the composition of 3D objects shown in a single image. Recent methods have shown promising composition editing results in the image setting, but in the video setting, editing methods have focused on editing object’s appearance and motion, or camera motion, and as a result, methods to edit object composition in videos are still missing. We propose VideoHandles as a method for editing 3D object compositions in videos of static scenes with camera motion. Our approach allows editing the 3D position of a 3D object across all frames of a video in a temporally consistent manner. This is achieved by lifting intermediate features of a generative model to a 3D reconstruction that is shared between all frames, editing the reconstruction, and projecting the features on the edited reconstruction back to each frame. To the best of our knowledge, this is the first generative approach to edit object compositions in videos. Our approach is simple and training-free, while outperforming state-of-the-art image editing baselines. Our project page is https://videohandles.github.io.
@@
**footnotetext: Work done during internship.
@@ 1 Introduction
Diffusion models and flow-based models are currently the standard for high-quality text-to-image generation. Text-to-video diffusion/flow-based models lag behind in quality, but have recently seen big improvements. The prevalent text-based control is easy to use, but impractical for some types of edits, such as edits of the object composition in a scene: specifying the position of an object with text is inaccurate and iterative editing workflows are not supported. Several recent methods address this issue in the image domain by proposing different types of iterative image editing methods. These either focus on editing the appearance of objects [47, 4, 13], or their spatial composition [29, 3, 1]. In the video domain, current methods support editing only the appearance [6, 23] while lacking methods to edit spatial object compositions, for example, editing the 3D position of objects in generated videos, as shown in Figure 1. Editing the object composition in a video introduces several challenges: a plausible editing output requires generating details such as shadows and lighting that may have changed due to the edited composition; furthermore, the edited video needs to preserve the identity of the original objects and should adhere to an edit control manipulated by the user. Finally, the edit needs to be applied to all video frames in a temporally consistent manner.
@@
We propose VideoHandles as a generative approach to edit the object composition in a video of a static scene. Our approach allows editing the 3D position of a 3D object in a video, resulting in a plausible, temporally consistent edit that preserves the identity of the original object. To the best of our knowledge, ours is the first generative approach that allows editing the object composition in a video. Given a pretrained flow-based video generative model, we present a novel method to edit the intermediate features from the generative model’s network in a temporally consistent manner. Specifically, we lift the intermediate features of each frame to a common 3D reconstruction, effectively treating them as latent textures. We then edit the 3D location of an object using 3D translations or rotations, and project the features back to their corresponding frames. We use such projected features as guidance during the generative process to create a plausible edited video. Editing of real (non-generated) videos is supported by first inverting them into the random noise. Our approach is simple and does not require any training or finetuning that risks biasing the distribution of the generative model.
@@
We evaluate our method on several generated and captured videos. As there are no existing methods that are specialized to editing the 3D object composition in videos, we compare to several image editing baselines that can be applied in a per-frame manner. We evaluate the results in terms of plausibility, temporal consistency, identity preservation, and adherence to the target edit. In addition to a large number of qualitative comparisons, we also conduct a user study. The users have a clear preference for our method in terms of plausibility and temporal consistency, while our method is at least on par or slightly better than image editing baselines in terms of identity preservation and edit adherence. Finally, we perform a quantitative evaluation which confirms these findings.
@@
  • For the feature-based generative editing process, we describe the optimal approach for feature extraction from the video generative model network (Section 4.2).
  • We also demonstrate that self-attention-map-based weighting (Section 4.4) and null-text prediction in the foreground region (Section 4.5) further improve the editing quality.
@@ 2 Related Work
There has been a series of work that focus on manipulating intermediate features or attention maps of pre-trained image diffusion models to edit the appearance of objects within an image [10, 42, 5, 47, 4, 13] in a zero-shot setting. While effective, such methods often do not focus on editing the composition of objects, which requires control over object positions and strict identity preservation. To tackle identity preservation, various customization approaches have been proposed that enable the generation of images of a particular object or subject in different compositions. However, such methods do not provide edit controls and typically require finetuning of the base model [33, 19]. The prior of image diffusion models has been further utilized to enable editing of 3D static scenes represented as 3D neural assets via iterative optimization approaches [30, 16, 18]. These methods, however, also focus on changing the appearance of objects, rather than our goal of composition editing.
@@
Several recent methods aim at editing the composition of objects in an image [26, 29, 9, 3, 8, 1, 49, 7, 52, 2]. Another line of work aims at inserting an object from a source image into a new target image [39, 40, 46], which can be repurposed as image editing tools by using the same image as source and target. Another popular editing workflow provides control points that can be dragged by a user to deform objects or edit 2D object positions [36, 28, 35, 22, 37]. Additionally, a few more general image editing methods have been proposed that can be used for either image appearance editing or image composition editing [51, 25]. All of these methods can be applied to videos by separately editing each frame, but this loses temporal consistency, as we show in our experiments in Section 5. Most related to our work is Diffusion Handles [29] which inspired our approach of editing intermediate features using a 3D reconstruction. We show how to modify this approach so it can be applied to non-depth-conditioned video priors, including which features to pick, which 3D reconstruction method to use, how to avoid artifacts from hard object masks, and how to effectively remove the original object from the edited video.
@@
With the increasing quality of video generators, various works have focused on editing the appearance of objects within videos. A key issue these works aim to tackle is to maintain temporal consistency between frames while changing the appearance. To address this, some works [6, 32, 7] have proposed techniques to maintain consistency using only image diffusion models, while others [15, 14] have leveraged priors from video diffusion models to tackle this challenge. While successful in preserving temporal consistency, these approaches are limited to appearance changes, and no prior work addresses changing object compositions in videos.
@@
Recently, another line of work in the video domain focuses on controlling the motion of objects or cameras during generation [43, 45, 34, 20]. Although these methods allow specifying how a particular object should move in a video, they are designed for generation rather than editing tasks, thus they do not allow modifying the compositions of static object arrangements in videos. Moreover, unlike these approaches that require training on task-specific datasets to learn motion control, VideoHandles is training-free.
@@ 4 VideoHandles: A 3D-Aware Video Editing Method
We first describe how to obtain a 3D-aware warping function in the 2D space of each frame. Given a set of 2D coordinates $\Omega_{H,W}=\{(v,u)\mid v\in[0,H),\,u\in[0,W)\}$, the connection between the 3D space and the $i$-th 2D frame is established through the projection function $f^{(i)}:\mathbb{R}^{3}\rightarrow\Omega_{H,W}$, which is defined by the $i$-th camera pose. Let $\mathcal{B}^{(1)}_{\text{src}}:\Omega_{H,W}\rightarrow\{0,1\}$ denote the 2D binary mask of an object selected by users in the first frame. Based on the 2D object mask in the first frame $\mathcal{B}^{(1)}_{\text{src}}$, we first partition $\boldsymbol{P}_{\text{src}}$, the point cloud reconstructed from the input video $\boldsymbol{X}_{\text{src}}$, as follows:
$\boldsymbol{P}_{f}$ =$=\{\mathbf{p}\in\boldsymbol{P}_{\text{src}}\mid\mathcal{B}^{(1)}_{\text{src}}(f^{(1)}(\mathbf{p}))=1\},$ (3)
@@
where $\boldsymbol{P}_{f}$ consists of points whose projections lie within the 2D masked region defined by $\mathcal{B}^{(1)}_{\text{src}}$, and $\boldsymbol{P}_{b}$ denotes the remaining points representing the background. By applying a 3D transformation $\mathcal{T}$ to $\boldsymbol{P}_{f}$ alone, we construct a rough target 3D scene represented as a point cloud:
@@ 5 Experiments
We also present snapshots of the edited videos in Figure 1 and Figure 4. Qualitatively, our method successfully edits object composition in videos while making appropriate contextual adjustments, such as the new reflection beneath the wine glass in Figure 1 and the new shadows beneath the transformed car, apple, and vase in rows 2, 3, and 5 of Figure 4, respectively. In comparison, Diffusion Handles [29] (the fourth column in Figure 4) alters the identity of objects or the background across different frames, as seen in the second row, and frequently duplicates objects, as shown in the first row. Direct frame warping (the second column) and its refined one by SDEdit [25] (the third column) also typically produce visual seams (second row) and implausible objects (fourth row) due to inaccuracies in warping.
@@
Please refer to the appendix for the edited video results. In the first row, the results without null-text (third column) exhibit object duplication, showing the armchair in both the original and target positions. In the second row, the results without self-attention-based weighting (second column) drastically alter the background colors, and the results without null-text (third column) introduce a new knob on the kettle. In contrast, our full method (last column) best preserves the identity of the kettle. In the third row, the results without self-attention-based weighting (second column) and with temporal layer features (fourth column) generate a new lamp next to the armchair and thus fail to preserve the background. Our method successfully moves the selected armchair without changing the background.
@@ 6 Conclusion
We have presented VideoHandles, the first method to our knowledge that leverages the prior of video generative models for editing object composition in video. Given the warping function for each frame obtained from a 3D reconstruction and transformation of an object in 3D space, VideoHandles applies temporally consistent warping to features extracted from a pre-trained video generative model, rather than to the frames themselves, using these features as guidance in the generative process. Experimental results, including a user study, demonstrate that VideoHandles outperforms per-frame editing methods in terms of plausibility, identity preservation, and edit coherence.
Limitations and Future Work.
@@ Acknowledgements
This work was supported by the NRF grant (RS-2023-00209723) and IITP grants (RS-2022-II220594, RS-2023-00227592, RS-2024-00399817), all funded by the Korean government (MSIT), as well as grants from the DRB-KAIST SketchTheFuture Research Center and NAVER-Intel.
@@ A.2 Temporal Consistency Across Frames
A.3 Inference Time and Computation Cost
As no other method attempts to edit object compositions in videos, we compare ours with the most technically related image editing method, DiffusionHandles [29]. While DiffusionHandles requires editing each frame individually, our method processes all frames in a single feed-forward pass, reducing the editing time to 6 minutes compared to DiffusionHandles’ 96 minutes for a 51-frame video on A6000 GPUs.
@@ A.4 Implementation Details
To define the object binary mask in the first frame, $\mathcal{B}_{\text{src}}^{(1)}$, we leverage SAM [17]. For the object transformation energy function, $\mathcal{G}_{o}$, we use features from all 24 spatial self-attention layers in OpenSora [54], while for the background preservation energy function, $\mathcal{G}_{b}$, we only use features from the first 14 self-attention layers. For both the input video generation and the guided generative process for editing, we use $30$ steps with a classifier-free-guidance scale of $7$. For the gradient step sizes of the energy functions discussed in Section 4.4, we set $\rho_{o}$ and $\rho_{b}$ to 650 and 100, respectively, for all editing examples in the comparisons. To better preserve the background details of the input video and smoothly adjust it to the new object composition, we initially compute the background preservation energy function without the averaging operator, measuring feature discrepancy element-wise. We then transition to the average loss to allow the background to adapt smoothly to the new object composition. This switch occurs at the ninth step out of 30 steps in the generative process.

### v2

@@ Abstract
Generative methods for image and video editing leverage generative models as priors to perform edits despite incomplete information, such as changing the composition of 3D objects depicted in a single image. Recent methods have shown promising composition editing results for images. However, in the video domain, editing methods have focused on editing objects’ appearance and motion, or camera motion. As a result, methods for editing object composition in videos remain largely unexplored. We propose VideoHandles as a method for editing 3D object compositions in videos of static scenes with camera motion. Our approach enables the editing of an object’s 3D position across all frames of a video in a temporally consistent manner. This is achieved by lifting intermediate features of a generative model to a 3D reconstruction that is shared between all frames, editing the reconstruction, and projecting the features on the edited reconstruction back to each frame. To the best of our knowledge, this is the first generative approach to edit object compositions in videos. Our approach is simple and training-free, while outperforming state-of-the-art image editing baselines. Our project page is https://videohandles.github.io.
@@
**footnotetext: Work done during an internship at Adobe Research.
@@ 1 Introduction
Diffusion and flow-based models are currently the standard for high-quality text-to-image generation. Text-to-video diffusion/flow-based models lag behind in quality, but have recently seen big improvements. The prevalent text-based control is easy to use, but impractical for some types of edits, such as edits of the object composition in a scene: specifying the position of an object with text is inaccurate and iterative editing workflows are not supported. Several recent methods address this issue in the image domain by proposing different types of iterative image editing methods. These either focus on editing the appearance of objects [47, 4, 13], or their spatial composition [29, 3, 1]. In the video domain, current methods support editing only the appearance [6, 23] while lacking methods to edit spatial object compositions, for example, editing the 3D position of objects in videos, as shown in Figure 1. Editing the object composition in a video introduces several challenges: a plausible editing output requires generating details such as shadows and lighting that may have changed due to the edited composition; furthermore, the edited video needs to preserve the identity of the original objects and should adhere to an edit control manipulated by the user. Finally, the edit needs to be applied to all video frames in a temporally consistent manner.
@@
We propose VideoHandles as a generative approach to edit the object composition in a video of a static scene. Our approach enables the editing of an object’s 3D position in a video, resulting in a plausible, temporally consistent edit that preserves the identity of the original object. To the best of our knowledge, ours is the first generative approach that allows editing the object composition in a video. Given a pretrained flow-based video generative model, we present a novel method to edit the intermediate features from the generative model’s network in a temporally consistent manner. Specifically, we lift the intermediate features of each frame to a common 3D reconstruction, effectively treating them as latent textures. We then edit the 3D location of an object using 3D translations or rotations, and project the features back to their corresponding frames. We use such projected features as guidance during the generative process to create a plausible edited video. Our approach is simple and does not require any training or finetuning that risks biasing the distribution of the generative model.
@@
We evaluate our method on several generated and captured videos. As there are no existing methods that are specialized to editing the 3D object composition in videos, we compare to several image editing baselines that can be applied in a per-frame manner. We evaluate the results in terms of plausibility, temporal consistency, identity preservation, and adherence to the target edit. In addition to a large number of qualitative comparisons, we also conduct a user study. The results show a clear preference for our method in terms of plausibility and temporal consistency, while our method is at least on par with, or slightly better than image editing baselines in identity preservation and edit adherence. Finally, a quantitative evaluation further supports these findings.
@@
  • We also demonstrate that self-attention-map-based weighting (Section 4.4) and null-text prediction in the foreground region (Section 4.5) further improve editing quality.
@@ 2 Related Work
There has been a series of work that focus on manipulating intermediate features or attention maps of pre-trained image diffusion models to edit the appearance of objects within an image [10, 42, 5, 47, 4, 13] in a zero-shot setting. While effective, such methods often do not focus on editing the composition of objects, which requires control over object positions and strict identity preservation. To tackle identity preservation, various customization approaches have been proposed that enable the generation of images of a particular object or subject in different compositions. However, such methods do not provide edit controls and typically require finetuning of the base model [33, 19]. The prior of image diffusion models has been further utilized to enable editing of 3D static scenes represented as 3D neural assets  [30, 16, 18]. These methods, however, also primarily focus on changing the appearance of objects, rather than our goal of composition editing.
@@
Several recent methods aim at editing the composition of objects in an image [26, 29, 9, 3, 8, 1, 49, 7, 52, 2]. Another line of work aims at inserting or removing objects in an image [39, 40, 46], where the combination of both can be considered as object composition editing. Another popular editing workflow provides control points that can be dragged by a user to deform objects or edit 2D object positions [36, 28, 35, 22, 37]. Additionally, a few more general image editing methods have been proposed that can be used for either image appearance editing or image composition editing [51, 25]. All of these methods can be applied to videos by separately editing each frame, but this loses temporal consistency, as we show in our experiments in Section 5. Most related to our work is DiffusionHandles [29] which inspired our approach of editing intermediate features using a 3D reconstruction. We show how to modify this approach so it can be applied to non-depth-conditioned video priors, including which features to pick, which 3D reconstruction method to use, how to avoid artifacts from hard object masks, and how to effectively remove the original object from the edited video.
@@
With the increasing quality of video generative models, various works have focused on editing the appearance of objects within videos. A key issue these works aim to tackle is to maintain temporal consistency between frames while changing the appearance. To address this, some works [6, 32, 7] have proposed techniques to maintain consistency using only image diffusion models, while others [15, 14] have leveraged priors from video diffusion models to tackle this challenge. While successful in preserving temporal consistency, these approaches are limited to appearance changes, and no prior work addresses changing object compositions in videos.
@@
Recently, another line of work in the video domain focuses on controlling the motion of objects or cameras during generation [43, 45, 34, 20]. Although these methods allow specifying how a particular object should move in a video, they are designed for generation rather than editing tasks, thus they do not allow modifying the compositions of object arrangements in existing videos. Moreover, unlike these approaches that require training on task-specific datasets to learn motion control, VideoHandles is training-free.
@@ 4 VideoHandles: A 3D-Aware Video Editing Method
We first describe how to obtain a 3D-aware warping function in the 2D space of each frame. Given a set of 2D coordinates $\Omega_{H,W}=\{(v,u)\mid v\in[0,H),\,u\in[0,W)\}$, the connection between the 3D space and the $i$-th 2D frame is established through the projection function $f^{(i)}:\mathbb{R}^{3}\rightarrow\Omega_{H,W}$, which is defined by the $i$-th camera pose. Let $\mathcal{B}^{(1)}:\Omega_{H,W}\rightarrow\{0,1\}$ denote the 2D binary mask of an object selected by users in the first frame. Based on the 2D object mask in the first frame $\mathcal{B}^{(1)}$, we first partition $\boldsymbol{P}_{\text{src}}$, the point cloud reconstructed from the input video $\boldsymbol{X}_{\text{src}}$, as follows:
$\boldsymbol{P}_{f}$ =$=\{\mathbf{p}\in\boldsymbol{P}_{\text{src}}\mid\mathcal{B}^{(1)}(f^{(1)}(\mathbf{p}))=1\},$ (3)
@@
where $\boldsymbol{P}_{f}$ consists of points whose projections lie within the 2D masked region defined by $\mathcal{B}^{(1)}$, and $\boldsymbol{P}_{b}$ denotes the remaining points representing the background. By applying a 3D transformation $\mathcal{T}$ to $\boldsymbol{P}_{f}$ alone, we construct a rough target 3D scene represented as a point cloud:
@@ 5 Experiments
Please refer to our project page***https://videohandles.github.io/ for the results presented in videos, including those with a more recent video generative model, CogVideoX [48], instead of OpenSora [54]. We also present snapshots of the edited videos in Figure 1 and Figure 4. Qualitatively, our method successfully edits object composition in videos while making appropriate contextual adjustments, such as the new reflection beneath the wine glass in Figure 1 and the new shadows beneath the transformed car, apple, and vase in rows 2, 3, and 5 of Figure 4, respectively. In comparison, DiffusionHandles [29] (the fourth column in Figure 4) alters the identity of objects or the background across different frames, as seen in the second row, and frequently duplicates objects, as shown in the first row. These failures are more evident in the videos shown in our project page. Direct frame warping (the second column) and its refined one by SDEdit [25] (the third column) also typically produce visual seams (second row) and implausible objects (fourth row) due to inaccuracies in warping.
@@
In the first row, the results without null-text (third column) exhibit object duplication, showing the armchair in both the original and target positions. In the second row, the results without self-attention-based weighting (second column) drastically alter the background colors, and the results without null-text (third column) introduce a new knob on the kettle. In contrast, our full method (last column) best preserves the identity of the kettle. In the third row, the results without self-attention-based weighting (second column) and with temporal layer features (fourth column) generate a new lamp next to the armchair and thus fail to preserve the background. Our method successfully moves the selected armchair without changing the background.
@@ 6 Conclusion
We have presented VideoHandles, the first method that leverages the prior of video generative models for editing object composition in videos. Given the warping function for each frame obtained from a 3D reconstruction and transformation of an object in 3D space, VideoHandles applies temporally consistent warping to features extracted from a pre-trained video generative model, rather than to the frames themselves, using these features as guidance in the generative process. Experimental results, including a user study, demonstrate that VideoHandles outperforms per-frame editing methods in terms of plausibility, identity preservation, and edit coherence.
Limitations and Future Work
@@ Acknowledgements
We appreciate Yunhong Min for his valuable assistance with our experiments. This work was supported by the NRF grant (RS-2023-00209723) and IITP grants (RS-2022-II220594, RS-2023-00227592, RS-2024-00399817), all funded by the Korean government (MSIT), as well as grants from the DRB-KAIST SketchTheFuture Research Center.
@@ A.2 Temporal Consistency Across Frames
A.3 Inference Time and Computational Cost
As no other method attempts to edit object compositions in videos, we compare ours with the most technically related image editing method, DiffusionHandles [29]. While DiffusionHandles requires editing each frame individually, our method processes all frames in a single feed-forward pass, reducing the editing time to 6 minutes compared to DiffusionHandles’ 96 minutes for a 51-frame video on A6000 GPUs. Please note that additional time may be accounted for due to inter-GPU communication overhead, as we split inference across two GPUs to accommodate VRAM constraints.
@@ A.4 Implementation Details
To define the object binary mask in the first frame, $\mathcal{B}^{(1)}$, we leverage SAM [17]. For the object transformation energy function, $\mathcal{G}_{o}$, we use features from all 24 spatial self-attention layers in OpenSora [54], while for the background preservation energy function, $\mathcal{G}_{b}$, we only use features from the first 14 self-attention layers. For both the input video generation and the guided generative process for editing, we use $30$ steps with a classifier-free-guidance scale of $7$. For the gradient step sizes of the energy functions discussed in Section 4.4, we set $\rho_{o}$ and $\rho_{b}$ to 650 and 100, respectively, for all editing examples in the comparisons. To better preserve the background details of the input video and smoothly adjust it to the new object composition, we initially compute the background preservation energy function without the averaging operator, measuring feature discrepancy element-wise. We then transition to the average loss to allow the background to adapt smoothly to the new object composition. This switch occurs at the ninth step out of 30 steps in the generative process.